OpenStudy (anonymous):

find the value of -p + 10 for each given value of p 1.p=1 2.p=2.55 3.p=10 4.p=11 neeed help :D

OpenStudy (anonymous):

1. -1 + 10 = 9 2. -2.55 + 10 = 7.45 3. -10 + 10 = 0 4. -11 + 10 = -1